I am supposed to write the equation of the line with x intercept (12,0) and y intercept (0,4). Can someone help me with this?
Answer by stanbon(48516)

I am supposed to write the equation of the line with x intercept (12,0) and y intercept (0,4).
-----------
You have two points:
Find the slope = [4-0]/[0-12) = -1/3
--------
The equation has the form y=mx+b where y=4 when x=0 and where m=-1/3
So, 4 = (-1/3)*0 + b
So, b=4
------------
EQUATION:
y=(-1/3)x+4
==================
Cheers,
Stan H.